OpenStudy (anonymous):

Select the equation of the line that passes through the point (–2, 1) and has slope 3 in point-slope form. A. (y – 1) = 3(x + 2) B. (x – 2) = 3(y + 1) C. (x + 2) = 3(y – 1) D. (y + 1) = 3(x – 2)

OpenStudy (anonymous):

So?

OpenStudy (***[isuru]***):

hey.. The standard point-slope equation looks like this: \[Y - Y _{1} = m( X - X _{1})\] It should be noted that "y1" or "x1" does not mean y multipled by 1. In this case it simply denotes a particular y value which you will plug into the equation. The variable m is the slope of the line if we take (-2 , 1) we can write X1 = -2 Y1 = 1 then -X1 = 2 -Y1 = -1 then the point slope form will be something like this...... \[Y - 1 = m( X +2 )\] and the gradient is 3 . so.... "m" will be 3 \[Y - 1 = 3( X +2 )\] now ... the rest is up to u :)

OpenStudy (anonymous):

So its A right?

OpenStudy (***[isuru]***):

yep!! :)